# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 6 Simulations
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.303 0.000 0.000 0.000 0.311 0.000 0.000 0.000
NCON 0.344 0.925 0.345 0.345 0.353 0.951 0.415 0.392
Index $0$ 0.724 0.986 $0$ 0.705 0.968
NW 2.672 0.695 53.569 0.316 3.952 0.881 104.417 0.377
rb-NW 0.409 0.266 1.913 0.047 0.472 0.343 3.063 0.047
POET 0.867 0.405 8.044 0.100 1.239 1.057 26.840 0.162
NLS 0.468 0.390 3.911 0.047 0.501 0.568 6.930 0.048
SFNL 0.429 0.264 2.337 0.044 0.489 0.337 3.519 0.073
  • See notes in Table 1.
We then turn to the settings with both TE and weight (equality) constraints. As described earlier, we impose weight restrictions on the first $10$ stocks. Here we set $\omega=0.2$. We keep the same DGP as the previous TE-constraint-only case and only change the way portfolios are formed due to the extra constraint. Tables 3 and 4 report the simulation results. We first notice that in Table 3, where $T=100$, rb-NW has the best performance in SR error for both $p<T$ and $p>T$, and all TE levels. The margin of SR error differences between rb-NW and other non-oracle portfolios are usually large. For example, at $p=120,T=100$, SR error at $TE=0.1$ for our method is 0.868, whereas the second best method SFNL has error of 1.037. It has similarly best performance in Tables 4, where $T=150$, except for $p=180$ and TE=0.2, where NLS has slightly better SR error performance (0.083 compared to 0.089). rb-NW again has good overall performance in terms of TE control, and is closest to the oracle level of risk, weight in most occasions. For example, in the case of Table 4 set $(T,p)=(150,120)$ and target TE = 0.1, rb-NW has TE at 0.116 which is close to the Oracle at 0.101. Even for the risk error, rb-NW is the best in Tables 3-4 among the comparison portfolios.
Next, we show simulation results for TE and weight inequality constraints. We use the same DGP and the restricted asset pool as in the TE and weight equality constraint, and set the parameters $\omega$ and $\kappa$ (discussed later) such that the weight inequality constraint is binding/non-binding in the population level. Intuitively, a good portfolio should behave similar to the TE and weight equality constraint optimal portfolio when the weight constraint is theoretically binding, and similar to the TE constraint only optimal portfolio when the weight constraint is nonbinding. Under our DGPs, with simple calculation, the smallest $\kappa w_{u}$ is -0.004, which is obtained when TE $=0.3$ and $p=120$, and the largest $\kappa w_{u}$ is 0.0104 which is obtained when TE $=0.3$ and $p=80$. We then set $\omega=0.004$ such that the weight constraints for four scenarios ($\{p=80,\text{TE}=0.2,0.3\}$, $\{p=180,\text{TE}=0.2,0.3\}$) are binding, and the rest scenarios are non-binding. Tables 5-6 show the results. Especially in the cases where $p>T$, it has the generally best performance in different metrics (uniformly best performance in SR in the right panels of Tables 5-6).
Table 3: Simulation results with TE and weight constraints, sample size $T=100$.
p=80, T=100 p=120, T=100
TE = 0.1
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.102 0.000 0.000 0.000 0.101 0.000 0.000 0.000
NCON 0.223 0.997 0.953 4.256 0.211 0.992 0.947 3.963
Index $0$ 0.775 0.330 $0$ 0.747 0.314
NW 0.339 0.803 2.409 2.848 0.512 1.349 6.241 3.454
rb-NW 0.110 0.371 0.198 0.532 0.115 0.580 0.405 0.868
POET 0.150 0.526 0.538 1.511 0.185 0.856 1.084 1.812
NLS 0.113 0.557 0.365 0.892 0.119 0.821 0.728 1.380
SFNL 0.113 0.379 0.266 0.690 0.118 0.570 0.517 1.037
TE = 0.2
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.202 0.000 0.000 0.000 0.204 0.000 0.000 0.000
NCON 0.229 0.936 0.704 0.697 0.249 0.984 0.666 0.878
Index $0$ 0.413 0.205 $0$ 0.436 0.093
NW 1.038 2.371 12.391 0.596 1.352 3.787 30.519 0.761
rb-NW 0.263 1.130 0.960 0.133 0.280 1.791 1.953 0.269
POET 0.505 1.481 2.713 0.408 0.527 2.442 5.212 0.556
NLS 0.303 1.477 1.714 0.241 0.324 2.316 3.575 0.419
SFNL 0.307 1.109 1.376 0.207 0.311 1.726 2.413 0.357
TE = 0.3
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.296 0.000 0.000 0.000 0.305 0.000 0.000 0.000
NCON 0.288 0.899 0.536 0.324 0.306 0.950 0.526 0.374
Index $0$ 0.271 0.174 $0$ 0.268 0.140
NW 1.475 3.456 13.546 0.262 2.681 6.104 40.793 0.329
rb-NW 0.343 1.686 1.313 0.041 0.418 2.748 2.511 0.056
POET 0.749 2.329 4.416 0.162 0.685 3.783 6.295 0.151
NLS 0.394 2.058 2.287 0.072 0.473 3.620 4.527 0.093
SFNL 0.395 1.663 1.911 0.059 0.465 2.643 3.118 0.083
  • “Oracle” is the portfolio using the population $\mu$ and $\Xi$ and plug in the analytical solution for $w_{cp}^{*}$ given the target TE and weight constraints. “NCON” is the oracle portfolio without TE or weight constraint. For other portfolios, see notes in Table 1, where the plug-in portfolios use the formula for $\hat{w}_{cp}$ instead of $\hat{w}_{d}$. “-” means the statistic is not available.
Table 4: Simulation results with TE and weight constraints, sample size $T=150$.
p=120, T=150 p=180, T=150
TE = 0.1
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.101 0.000 0.000 0.000 0.100 0.000 0.000 0.000
NCON 0.214 0.985 0.954 2.742 0.193 1.005 0.922 13.634
Index $0$ 0.729 0.500 $0$ 0.777 0.008
NW 0.463 1.075 3.733 1.939 0.761 1.622 9.224 12.127
rb-NW 0.116 0.437 0.226 0.364 0.113 0.696 0.378 2.667
POET 0.189 0.712 0.734 1.176 0.194 1.383 1.958 6.432
NLS 0.117 0.701 0.394 0.603 0.117 1.177 0.878 4.245
SFNL 0.118 0.438 0.310 0.443 0.114 0.692 0.440 3.049
TE = 0.2
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.203 0.000 0.000 0.000 0.205 0.000 0.000 0.000
NCON 0.222 0.947 0.679 0.610 0.199 0.939 0.687 0.445
Index $0$ 0.383 0.204 $0$ 0.340 0.318
NW 1.503 3.991 30.951 0.547 2.383 5.795 55.584 0.421
rb-NW 0.253 1.439 1.110 0.118 0.271 2.215 1.942 0.089
POET 0.504 2.143 4.003 0.249 0.657 11.866 123.131 0.201
NLS 0.287 2.136 2.133 0.168 0.289 3.633 3.968 0.083
SFNL 0.308 1.418 1.452 0.152 0.284 2.172 2.396 0.114
TE = 0.3
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.302 0.000 0.000 0.000 0.308 0.000 0.000 0.000
NCON 0.331 0.881 0.604 0.484 0.289 0.933 0.511 0.272
Index $0$ 0.355 0.147 $0$ 0.228 0.185
NW 2.147 5.238 21.557 0.448 3.281 9.097 81.857 0.240
rb-NW 0.396 2.220 1.305 0.072 0.454 3.462 2.638 0.028
POET 0.874 3.449 5.023 0.324 1.081 7.100 10.094 0.084
NLS 0.459 3.073 2.452 0.139 0.470 5.494 5.503 0.048
SFNL 0.438 2.209 1.632 0.111 0.465 3.377 3.059 0.028
  • See notes in Table 3.
Table 5: Simulation results with TE constraint and inequality constraint, $\omega=0.004$, $T=100$.
p=80, T=100 p=120, T=100
TE = 0.1, All Nonbinding
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.100 0.000 0.000 0.000 0.102 0.000 0.000 0.000
NCON 0.255 0.987 0.863 4.030 0.212 0.975 0.802 2.531
Index $0$ 0.816 0.035 $0$ 0.739 0.037
NW 0.529 1.210 4.296 3.274 0.904 2.208 17.007 2.446
rb-NW 0.132 0.605 0.388 0.770 0.159 0.961 0.992 0.867
POET 0.189 0.681 0.750 1.214 0.269 1.147 2.036 1.069
NLS 0.167 0.743 0.747 1.409 0.181 1.325 1.937 1.232
SFNL 0.156 0.596 0.548 1.097 0.177 0.926 1.216 1.015
TE = 0.2, Binding with $p=80,T=100$, Nonbinding $p=120,T=100$
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.199 0.000 0.000 0.000 0.203 0.000 0.000 0.000
NCON 0.301 0.973 0.623 1.189 0.268 0.950 0.523 0.737
Index $0$ 0.559 0.015 $0$ 0.447 0.018
NW 1.074 2.419 11.567 1.075 1.836 4.416 41.448 0.702
rb-NW 0.274 1.209 1.035 0.294 0.311 1.922 2.352 0.214
POET 0.437 1.354 1.953 0.446 0.530 2.285 4.646 0.260
NLS 0.331 1.485 2.006 0.542 0.355 2.651 4.619 0.304
SFNL 0.315 1.192 1.442 0.446 0.345 1.853 2.833 0.285
TE = 0.3, Binding with $p=80,T=100$, Nonbinding $p=120,T=100$
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.297 0.000 0.000 0.000 0.303 0.000 0.000 0.000
NCON 0.368 0.960 0.439 0.563 0.344 0.925 0.345 0.345
Index $0$ 0.375 0.011 $0$ 0.276 0.014
NW 1.602 3.607 17.128 0.512 2.738 6.583 57.007 0.313
rb-NW 0.409 1.803 1.528 0.086 0.463 2.865 3.206 0.040
POET 0.651 2.013 2.831 0.159 0.697 3.397 6.182 0.066
NLS 0.493 2.213 2.969 0.211 0.529 3.952 6.308 0.061
SFNL 0.469 1.777 2.118 0.335 0.515 2.762 3.838 0.062
Table 6: Simulation results with TE constraint and inequality constraint, $\omega=0.004$, $T=150$.
p=120, T=150 p=180, T=150
TE = 0.1, All Non-binding
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.102 0.000 0.000 0.000 0.103 0.000 0.000 0.000
NCON 0.212 0.975 0.802 2.531 0.229 0.984 0.838 2.522
Index $0$ 0.739 0.037 $0$ 0.739 0.082
NW 0.896 2.137 16.004 2.442 1.322 3.331 29.371 2.431
rb-NW 0.137 0.774 0.594 0.614 0.158 1.206 0.908 0.814
POET 0.193 1.027 1.884 0.943 0.202 2.361 4.053 1.004
NLS 0.157 1.134 1.201 0.946 0.168 1.996 2.004 1.099
SFNL 0.144 0.761 0.750 0.733 0.164 1.173 1.059 0.920
TE = 0.2 Nonbinding with $p=120,T=150$, Binding $p=180,T=150$
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.203 0.000 0.000 0.000 0.207 0.000 0.000 0.000
NCON 0.268 0.950 0.523 0.737 0.278 0.968 0.592 0.814
Index $0$ 0.447 0.018 $0$ 0.466 0.042
NW 1.792 4.274 38.960 0.704 2.645 6.662 72.796 0.791
rb-NW 0.274 1.547 1.400 0.153 0.316 2.411 2.164 0.243
POET 0.348 2.013 4.000 0.224 0.405 4.724 10.011 0.292
NLS 0.314 2.268 2.855 0.221 0.336 3.993 4.865 0.267
SFNL 0.288 1.522 1.726 0.205 0.328 2.345 2.497 0.296
TE = 0.3, Nonbinding with $p=120,T=150$, Binding $p=180,T=150$
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.303 0.000 0.000 0.000 0.308 0.000 0.000 0.000
NCON 0.344 0.925 0.345 0.345 0.351 0.952 0.419 0.400
Index $0$ 0.276 0.014 $0$ 0.299 0.033
NW 2.672 6.372 53.566 0.316 3.944 9.932 103.066 0.385
rb-NW 0.409 2.306 1.904 0.045 0.472 3.595 3.024 0.048
POET 0.519 3.001 5.368 0.081 0.604 7.043 14.156 0.123
NLS 0.468 3.382 3.894 0.047 0.500 5.952 6.839 0.049
SFNL 0.429 2.270 2.327 0.046 0.488 3.496 3.475 0.077
@@ 7 Empirical study
In the empirical study, we use monthly return data of 254 stocks enlisted in S&P 500 index with complete observations from 01/1981-12/2020. We first consider the constrained portfolio with TE constraint only. The following two scenarios of portfolio composition are considered: (1) Largest 100 stocks measured by market values. The out of sample period is from 01/1991 to 12/2020, which contains 360 periods, and the window size is set to 120; (2) All 254 stocks, the out of sample period is 01/2006 - 12/2020, which contains 180 periods, the window size is set to 180. Scenarios (1) and (2) represent the cases that sample size $T$ is larger (less) than portfolio size $p$, respectively. The portfolio is rebalanced monthly. Our forecasts are rolling-window setup. Assume that our first window is $1\mathrel{\mathop{:}}n_{I}$, in Scenario 1 for example, $n_{I}=120$. After weights are estimated in the window, as well as estimates for $\mu,\Sigma_{y}$ and then we report metrics using these three estimates for the next period $n_{I+1}$. In case of Scenario 1, after getting estimates for the first 120 months, we use then to calculate metrics like Sharpe-Ratio for 121 month in the sample as out-sample forecast. Then we repeat the same process by ignoring the first period in-sample, but with the same window length. In other words, in case of Scenario 1, we use $\mathrel{\mathop{:}}12$ 121 periods to have a forecast for $122$ period. We repeat this process with fixed window length until the end of out-sample and we average all out-sample, and report in Tables.

### v2

@@ 6 Simulations
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.3000 0.0000 0.0000 0.0000 0.0354 0.3000 0.0000 0.0000 0.0000 0.0354
NCON 0.3988 2.3658 0.4532 0.0616 0.0365 0.3975 2.3728 0.4531 0.0618 0.0365
Index - 1.9099 0.0800 0.0777 0.0340 - 1.9099 0.0766 0.0481 0.0346
NW 1.9454 7.4410 2.9154 0.1940 0.0318 1.9983 7.9864 2.9966 0.1628 0.0324
CROWN 0.2987 0.6426 0.0020 0.0012 0.0354 0.2999 0.4710 0.0019 0.0008 0.0354
POET 0.3128 8.4775 10.5361 0.0480 0.0342 0.3568 5.2820 10.4017 0.0275 0.0348
NLS 0.3265 8.1355 0.9194 0.0644 0.0343 0.3626 6.4713 0.4374 0.0449 0.0346
SFNL 0.4243 4.6117 0.5197 0.0343 0.0348 0.3707 3.2449 0.2373 0.0181 0.0351
@@ 7 Empirical Study
In the empirical study, we show the performance of the proposed portfolio using monthly return data of S&P 500 index component stocks. We consider the portfolio with 1) TE constraint only and 2) joint TE and weight inequality constraints, which are two leading examples in the mutual fund industry. The following three tracked benchmark indices are considered: (1) S&P 500 index. (2) S&P 500 Health Care index and (3) S&P 500 Energy index. Notice all three indices’ component stocks change over time. To avoid the look-ahead survival bias, we follow the approach of Barroso and Saxena (2022) to obtain portfolio weight estimates. When constructing the portfolios, we consider all available S&P 500 constituent stocks at each point of the out-of-sample period. In the Appendix, we also consider the S&P 500 Industrials index as the benchmark, as well as the characteristics and industry-sorted portfolios. The portfolio is rebalanced monthly.
Table 5 analyzes the S&P 500 index with TE constraints, Tables 6-7 analyze joint weight and TE inequality constraints in the S&P 500 Healthcare index and S&P 500 Energy index, respectively. Our forecasts are in a rolling-window setup, and the window size is set to 180. The full sample period for S&P 500 index study is 01/1981 to 12/2020, and the out-of-sample period is 01/1996 to 12/2020. The entire sample period for the two themed indices (healthcare and energy) is 01/1991 to 12/2020, and the out-of-sample period is 01/2006 to 12/2020, due to the availability of themed index under consideration. The in-sample period is from $1:T_{I}$, and the out-sample period is from $T_{I+1}:T$. Our first window is $1:T_{I}$, where $T_{I}=180$. After getting estimates for the first 180 months, we use them to calculate metrics like Sharpe Ratio for the 181st month in the sample as out-of-sample forecast. Then we repeat the same process by ignoring the first period in-sample but with the same window length. We repeat this process with a fixed window length until the end of out-sample, and we average all out-sample and report in the empirical results tables the average returns (AVR), TE (TE), standard deviation (Risk), and Sharpe Ratio (SR). We follow Caner et al. (2023) to obtain out-of-sample portfolio return and variance of the portfolio. After weights are estimated in the window, and defining the weight estimate at the first window as $\hat{\boldsymbol{w}}_{T_{I}}$, the out-of-sample portfolio return $\boldsymbol{w}^{\prime}\boldsymbol{\mu}$ is estimated by $(\widehat{\boldsymbol{w}^{\prime}\boldsymbol{\mu}})_{oos}:=\frac{1}{T-T_{I}}\sum_{t=T_{I}}^{T-1}\hat{\boldsymbol{w}}_{t}^{\prime}\boldsymbol{y}_{t+1}$, and for example $\hat{\boldsymbol{w}}_{t}$ at $t=T_{I}$ is the weight estimate in the window $1:T_{I}$. The out-of-sample variance of the portfolio, $\boldsymbol{w}^{\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}$ is estimated by $\frac{1}{T-T_{I}-1}\sum_{t=T_{I}}^{T-1}[\hat{\boldsymbol{w}}_{t}^{\prime}\boldsymbol{y}_{t+1}-(\widehat{\boldsymbol{w}^{\prime}\boldsymbol{\mu}})_{oos}]^{2}.$